OpenStudy (anonymous):

Solve the quadratic function x2 − 6x + 9 = 0 for x.

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

If you factor this you will get (x-3)(x-3) Use the zero property rule x-3 =0 x = 0+3 x=3